Example of a paper for LREC 2008

Author1, Author2, Author3

Affiliation1, Affiliation2, Affiliation3 Address1, Address2, Address3 author1@xxx.yy, author2@zzz.edu, author3@hhh.com

Abstract

Here is where you should put the abstract. Here is where you should put the abstract. Here is where you should put the abstract. Here is where you should put the abstract. Here is where you should put the abstract. Here is where you should put the abstract. Here is where you should put the abstract.

1. Introduction

Main body text.

2. Goal of the paper

Description of the main goal of the paper. Description of the main goal of the paper. Description of the main goal of the paper. Description of the main goal of the paper. Description of the main goal of the paper. Description of the main goal of the paper. Description of the main goal of the paper. Description of the main goal of the paper.

2.1. Example of a subsection An example of a subsection. An example of a subsection. An example of a subsection. An example of a subsection. An example of a subsection. An example of a subsection. An example of a subsection. An example of a subsection. An example of a subsection. An example of a subsection.

2.1.1. Example of a sub-subsection Yet another example, this time of a sub-subsection. Yet another example, this time of a sub-subsection. Yet another example, this time of a sub-subsection. Yet another example, this time of a sub-subsection. Yet another example, this time of a sub-subsection. Yet another example, this time of a sub-subsection. Yet another example, this time of a sub-subsection. Yet another example, this time of a sub-subsection. Yet another example, this time of a sub-subsection. Yet another example, this time of a sub-subsection. Yet another example, this time of a sub-subsection.

2.1.2. Example of a sub-subsection with a long

heading that will occupy two lines Yet another example of a sub-subsection. Yet another example of a sub-subsection. Yet another example of a subsubsection. Yet another example of a sub-subsection. Yet another example of a sub-subsection.

3. Additional guidelines

3.1. Footnotes This is an example of a footnote^{1}.

3.2. Figures

Example of a figure enclosed in a box.

This is a figure with a caption. This is a figure with a caption.

Figure 1: The caption of the figure.

3.3. Tables

Two types of tables are distinguished: in-column and big tables that don’t fit in the columns.

3.4. In-column tables

An example of an in-column table is presented here.

Level  Tools
Morphology Pitrat Analyser
Syntax  LFG Analyser (C-Structure)
Semantics  LFG F-Structures + Sowa’s

Conceptual Graphs

Table 1: The caption of the table

3.5. Big tables

An example of a big table which extends beyond the column and will float in the next page.

4. Citation Format

All references within the text are placed in parentheses containing the author’s surname followed by a comma before the date of publication (Chomsky, 1973). If the sentence already includes the author’s name, then it is only necessary to put the date in parentheses: Aslin (1949) . When several works are cited, those references are separated with a semicolon (Chomsky, 1973; Feigl, 1958): When the reference has more than three authors, only the name of the first author, followed by et al, appears (Fletcher et al., 1907). Bibliographical references are listed in alphabetical order at the end of the article. The title of the section, “References”, will be a level 1 heading. The first line of each bibliographical reference is justified to the left of the column, and the

1This is an example of the footnote text.

Level  Tools
Morphology Pitrat Analyser
Syntax  LFG Analyser (C-Structure)
Semantics  LFG F-Structures + Sowa’s Conceptual Graphs

Table 2: The caption of the big table

rest of the entry is indented by 0.35 cm. The following examples illustrate the format for conference proceedings (Chave, 1964), books (Butcher, 1981), articles (Howells, 1951) in journals, Ph.D. theses (Croft, 1978), and chapters of books (Feigl, 1958).

5. References

E. J. Aslin. 1949. Photostat recording in library work.

Aslib Proceedings, 1:49–52. Judith Butcher. 1981. Copy-editing. Cambridge Univer-

sity Press, 2nd edition. K. E. Chave. 1964. Skeletal durability and preservation. In

J. Imbrie and N. Newel, editors, Approaches to paleoecology, pages 377–87, New York. Wiley.

N. Chomsky. 1973. Conditions on transformations. In S. R. Anderson and P. Kiparsky, editors, A festschrift for Morris Halle, New York. Holt, Rinehart & Winston.

W. B. Croft. 1978. Organizing and searching large files of document descriptions. Ph.D. thesis, Cambridge University.

F. Feigl, 1958. Spot Tests in Organic Analysis, chapter 6.

Publisher publisher, 5th edition. W. M. Fletcher, F. G. Hopkins, M. Flet, and F. G. Hop.

1907. Lactic acid in amphibian muscle. J. Physiol., 35:247–309.

W. W. Howells. 1951. Factors of human physique. American Journal of Physical Anthropology, 9:159–192.